# Fast online node labeling with graph subsampling

### v1

@@ 1 Introduction
Large data applications like search and recommendation systems, rely on data stored in the form of very large, sparse, irregular graphs, where the number of nodes can be on the order millions, billions, or even trillions. In such cases, graph-based methods, such as node prediction, must accomplish their tasks exclusively using local operations, e.g. where the memory complexity is independent of graph size. This is the intention of methods like the localized approximate Personalized Page Rank method (APPR) (Andersen et al., 2006; Page et al., 1999), which approximates solving a sparse linear system by truncating messages whenever the residual of that coordinate is small. However, the complexity for these methods are often in terms of the maximum node degree, and their benefits are tied to the assumption that the graph node degrees are relatively uniform.
In this work, we consider a simple solution to this problem, where high-degree nodes subsample their neighbors in message-passing methods. This solution can save considerable memory overhead when the graph’s node degree distribution is heavily skewed. (See, for example, Figure 1.) However, the disadvantage of this strategy is that the stochasticity leads a high variance between each trial; thus, we implement present a mechanism for grounding the residual at each iteration, to reduce this variance.
@@
In this paper, we extend the work of (Andersen et al., 2006; Zhou et al., 2023) to graphs with unfavorable node degrees, using edge subsampling. We propose a simple approach: for a threshold $\bar{q}$, we identify all nodes with degree exceeding $\bar{q}$, and subsample their neighboring edges until they have $\leq\bar{q}$ neighbors. The remaining edges are reweighted such that the expected edge weights are held consistent. In offline graphs, sparsifications of this kind is $O(n)$ where $n$ is the number of nodes; however, the dependency on $n$ is removed when done online. Our contributions are
@@ 1.1 Related works
This problem category has been investigated by Belkin et al. (2004); Zhu et al. (2003); Herbster & Pontil (2006); Rakhlin & Sridharan (2017), and many others. Using ${\bm{z}}_{t}$ as a vector containing the already revealed label, then the solution to the APPR system is an approximation of those offered in Belkin et al. (2004); Zhu et al. (2003), for choices of regularization and interpolation. Relatedly, node prediction via approximate Laplacian inverse is also related to mean field estimation using truncated discounted random walks (Li et al., 2019), with known performance guarantees. The series of papers (Herbster & Pontil, 2006; Herbster et al., 2005) directly attack the suboptimality in online learning when applied to graphs with large diameters. They show that for this class of problems, using direct interpolation, the worst-case rate cannot be sublinear, and suggest additional graph structures to assist in this regime. Finally, Rakhlin & Sridharan (2017) tackled the problem of computing an online learning bound by offering a method, which can be seen as a modifined right-hand-side of the linear system in Belkin & Niyogi (2003). The advantage of Rakhlin & Sridharan (2017) is that it provides a means of computing a learning regret bound. In Zhou et al. (2023), the analysis was tightened to $O(\sqrt{n})$ sublinear regret bounds, under the appropriate choice of kernel.
@@ 2 Using graphs for node labeling
$\hat{y}_{t}=\operatorname{sign}\left(\frac{\sum_{i\in{\mathcal{N}}(t)}{\bm{A}}_{i,t}\tilde{y}_{i}}{\sum_{j\in{\mathcal{N}}(t)}{\bm{A}}_{j,t}}\right)=\operatorname{sign}\left({\bm{e}}_{t}^{T}{\bm{A}}{\bm{D}}^{-1}\tilde{\bm{y}}\right).$
@@
$\hat{y}_{t,\mathrm{soft}}=\sum_{k=0}^{K}\sum_{i\in{\mathcal{N}}(t)}(\beta^{k-1}{\bm{T}}^{k})_{i,t}\tilde{y}_{i}=\beta^{-1}\sum_{k=0}^{K}{\bm{e}}_{t}^{T}(\beta{\bm{A}}{\bm{D}}^{-1})^{k}\tilde{\bm{y}}\overset{K\to\infty}{=}\beta^{-1}{\bm{e}}_{t}^{T}({\bm{I}}-\beta{\bm{A}}{\bm{D}}^{-1})^{-1}\tilde{\bm{y}}.$ (1)
@@ 2.1 Approximate Personalized Page Rank
The APPR method (Andersen et al., 2006), (Alg. 1) is a memory-preserving approximation of PPR. Specifically, by having proper termination steps in the Push substep, the method offers a better tradeoff between learnability and memory locality. It solves the linear system equivalent to that of (PPR-symm), with $\beta=\frac{1-\alpha}{1+\alpha}\in(0,1)$ as the teleportation parameter.
@@
The following Lemma from Andersen et al. (2006) form the basis of the memory locality guarantee of this method.
@@ 3 Influencer-targeted sparsification
Figure 2: Is it better to subsample infleuncers? Correlation of inverse resistive distance (conductance, spectrally optimal) with high degree connecting edge is largely positive, across several graphs.
@@ 3.1 Resistive distance
Figure 3 investigates the performance of edge subsampling as it pertains to node labeling, our desired downstream task. This suggests that influencer-based sparsity not only preserves but can sometimes improve the edge ratio despite a reduction in the number of edges (even moreso than resistive subsampling).
@@ 3.2 Graph sparsification
Algorithm 2 General offline sparsifier 1:Graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},\mathbf{A})$, 2:edge probabilities $p_{u,v}$ for $(u,v)\in{\bm{E}}$ 3:Initialize $\tilde{\mathcal{E}}=\emptyset$, $\tilde{{\bm{A}}}=\mathbf{0}$, $\tilde{\mathcal{G}}=({\mathcal{V}},\tilde{\mathcal{E}},\tilde{{\bm{A}}})$ 4:for $i=1,...,m$ do 5:     Randomly pick edge $(u,v)$. 6:     With probability $p_{u,v}$, 7:$\tilde{{\bm{A}}}_{u,v}=\tilde{{\bm{A}}}_{v,u}=\frac{1}{p_{u,v}},\;\tilde{\mathcal{E}}=\tilde{\mathcal{E}}\bigcup\{(u,v)\}$ return $\tilde{\mathcal{G}}=({\mathcal{V}},\tilde{\mathcal{E}},\tilde{\mathbf{A}})$ [Uncaptioned image] Figure 3: Edge ratio. This measures the proportion of edges that connect different-labeled nodes over same-labeled nodes. (Smaller is better.) Sparsifications: U = uniform, R = resistive, I = influencer. The labels show $\bar{q}$ for (I), and the corresponding sparsification rate for (U) and (R).
@@
Consider ${\bm{L}}={\bm{I}}-{\bm{D}}^{-1/2}{\bm{A}}{\bm{D}}^{-1/2}$, $\tilde{\bm{L}}$ the Laplacian matrices corresponding to a graph and its sparsified version, and $\bar{q}$ the subsampling threshold. Then, for any ${\bm{x}}\in\mathbb{R}^{n}$, the term ${\bm{x}}^{T}\tilde{{\bm{L}}}_{I}{\bm{x}}$ is subgaussian with parameter
${\mathbf{subgauss}}({\bm{x}}^{T}\tilde{{\bm{L}}}_{I}{\bm{x}})=\frac{1}{4}\frac{d_{\max}^{2}}{\bar{q}^{2}}\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|.\vspace{-2ex}$
As a consequence,
$\mathbf{Pr}({\bm{x}}^{T}{\tilde{\bm{L}}_{I}}{\bm{x}}-{\bm{x}}^{T}{{\bm{L}}}{\bm{x}}\geq\epsilon)\leq\exp\left(-\frac{\epsilon^{2}\bar{q}^{2}}{d_{\max}^{2}\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|}\right).$
For a fixed probability, $\bar{q}=O(1/\epsilon)$. In total, the sampling complexity is $O(n/\epsilon)$. In comparison, optimal samplers have complexity $O(1/\epsilon^{2})$ (Karger, 1994; Benczúr & Karger, 2015; Spielman & Srivastava, 2008). For $n=O(1/\epsilon)$, the two rates are thus comparable. This lemma is most meaningful when $d_{\max}\gg\bar{q}$. For example, this is the regime if we set $\bar{q}$ to be a small multiple of the median node degree (see Table 1).
@@ 5.2 Unsupervised clustering
$\mathbf{cluster}(u)=\textsc{PickOne}(\underset{j\in{\mathcal{C}}}{\mathrm{arg\,max}}\{{\bm{Z}}_{t,j}\})$
@@ 6 Numerical experiments
Figure 6 gives the performance of offline and online (with and without dual correction) methods for online node labeling. Note that for small graphs, the baseline (WMA) is very strong; however, for larger graphs, the tradeoff for reduced misclasification rate becomes more apparent. It is also clear that the Relaxation method, although providing strong rates, is not as strong in practice as the Regularization method. Nonetheless, both methods are closely related, and their respective advantages and disadvantages stem from subtle adjustments in procedures and hyperparameters.
@@ B.1 General facts about subgaussianity
${\mathbf{subgauss}}(X)=S(p):=\begin{cases}1/4&1/4<S_{1}(p),0\leq p\leq 1\\ S_{1}(p)&1/4\geq S_{1}(p),0\leq p\leq 1\\ gg1&p>1\\ \end{cases}$
@@ B.2 Offline sparsification
${\mathbf{subgauss}}({\bm{x}}^{T}\tilde{{\bm{L}}}_{I}{\bm{x}})=\frac{d_{\max}^{2}}{4\bar{q}^{2}}\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|$
@@
Define $\tilde{{\bm{L}}}_{I}=\tilde{{\bm{L}}}-{\bm{L}}_{C}$.
@@
and $c_{i,j}$ for $(i,j)\not\in{\mathcal{S}}_{I}$. Then $\mathbb{E}[\tilde{{\bm{L}}}]={\bm{L}}$ and ${\bm{x}}^{T}\tilde{{\bm{L}}}{\bm{x}}$ is subgaussian with parameter $\sigma^{2}$ if and only if ${\bm{x}}^{T}\tilde{{\bm{L}}_{I}}{\bm{x}}$ is subgaussian with parameter $\sigma^{2}$. Since $c_{i,j}$ is a scaled Bernoulli random variable which is subgaussian with parameter $\sigma_{i,j}^{2}=\frac{S(p_{i,j})}{p_{i,j}^{2}}=S(\frac{\bar{q}}{\max\{d_{i},d_{j}\}})\frac{\max\{d_{i},d_{j}\}^{2}}{\bar{q}^{2}}$ and
@@
$\sum_{(i,j)\in{\mathcal{S}}_{I}}({\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j})^{2}b_{i,j}^{2}=\sum_{(i,j)\in{\mathcal{S}}_{I}}({\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j})^{2}S(\frac{\bar{q}}{\max\{d_{i},d_{j}\}})\frac{\max\{d_{i},d_{i}\}^{2}}{\bar{q}^{2}}\\ \leq S(\frac{\bar{q}}{\max\{d_{i},d_{j}\}})\frac{\max\{d_{i},d_{j}\}^{2}}{\bar{q}^{2}}\sum_{(i,j)\in{\mathcal{S}}_{I}}({\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j})^{2}.$
@@
${\mathbf{subgauss}}({\bm{x}}^{T}\tilde{{\bm{L}}}_{I}{\bm{x}})=\frac{d_{\max}^{2}}{\bar{q}^{2}}\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|$

### v2

@@ 1 Introduction
Large data applications like search and recommendation systems, rely on data stored in the form of very large, sparse, irregular graphs, where the number of nodes can be on the order millions, billions, or even trillions. In such cases, graph-based methods, such as node prediction, must accomplish their tasks exclusively using local operations, e.g. where the memory complexity is independent of graph size. This is the intention of methods like the localized approximate Personalized Page Rank method (APPR) (Andersen et al., 2006; Page et al., 1999), which approximates solving a sparse linear system by truncating messages whenever the residual of that coordinate is small. However, the complexity for these methods is often in terms of the maximum node degree, and their benefits are tied to the assumption that the graph node degrees are relatively uniform.
In this work, we consider a simple solution to this problem, where high-degree nodes subsample their neighbors in message-passing methods. This solution can save considerable memory overhead when the graph’s node degree distribution is heavily skewed. (See, for example, Figure 1.) However, the disadvantage of this strategy is that the stochasticity leads a high variance between each trial; thus, we implement a mechanism for grounding the residual at each iteration, to reduce this variance.
@@
In this paper, we extend the work of (Andersen et al., 2006; Zhou et al., 2023) to graphs with unfavorable node degrees, using edge subsampling. We propose a simple approach: for a threshold $\bar{q}$, we identify all nodes with degree exceeding $\bar{q}$, and subsample their neighboring edges until they have $\leq\bar{q}$ neighbors. The remaining edges are reweighted such that the expected edge weights are held consistent. In offline graphs, sparsifications of this kind require at least one full sweep through the graph, and grows with $n$, where $n$ is the number of nodes. However, online sparsification reduces this dependency on $n$, depending solely on the neighborhood structure of the query node. Our contributions are
@@ 1.1 Related works
This problem category has been investigated by Belkin et al. (2004); Zhu et al. (2003); Herbster & Pontil (2006); Rakhlin & Sridharan (2017), and many others. Using ${\bm{z}}_{t}$ as a vector containing the already revealed label, then the solution to the APPR system is an approximation of those offered in Belkin et al. (2004); Zhu et al. (2003), for choices of regularization and interpolation. Relatedly, node prediction via approximate Laplacian inverse is also related to mean field estimation using truncated discounted random walks (Li et al., 2019), with known performance guarantees. The series of papers (Herbster & Pontil, 2006; Herbster et al., 2005) directly attack the suboptimality in online learning when applied to graphs with large diameters. They show that for this class of problems, using direct interpolation, the worst-case rate cannot be sublinear, and suggest additional graph structures to assist in this regime. Finally, Rakhlin & Sridharan (2017) tackled the problem of computing an online learning bound by offering a method, which can be seen as a modified right-hand-side of the linear system in Belkin & Niyogi (2003). The advantage of Rakhlin & Sridharan (2017) is that it provides a means of computing a learning regret bound. In Zhou et al. (2023), the analysis was tightened to $O(\sqrt{n})$ sublinear regret bounds, under the appropriate choice of kernel.
@@ 2 Using graphs for node labeling
$\hat{y}_{t}=\operatorname{sign}\left(\frac{\sum_{i\in{\mathcal{N}}(t)}{\bm{A}}_{i,t}\tilde{y}_{i}}{\sum_{j\in{\mathcal{N}}(t)}{\bm{A}}_{j,t}}\right)=\operatorname{sign}\left({\bm{e}}_{t}^{T}{\bm{A}}{\bm{D}}^{-1}\tilde{\bm{y}}\right).\vspace{-2ex}$
@@
$\hat{y}_{t,\mathrm{soft}}=\sum_{k=0}^{K}\sum_{i\in{\mathcal{N}}(t)}(\beta^{k-1}{\bm{T}}^{k})_{i,t}\tilde{y}_{i}=\beta^{-1}\sum_{k=0}^{K}{\bm{e}}_{t}^{T}(\beta{\bm{A}}{\bm{D}}^{-1})^{k}\tilde{\bm{y}}\overset{K\to\infty}{=}\beta^{-1}{\bm{e}}_{t}^{T}({\bm{I}}-\beta{\bm{A}}{\bm{D}}^{-1})^{-1}\tilde{\bm{y}}.\vspace{-2ex}$ (1)
@@ 2.1 Approximate Personalized Page Rank
The APPR method (Andersen et al., 2006), (Alg. 1) is a memory-preserving approximation of PPR. Specifically, by having proper termination steps in the Push substep, the method offers a better tradeoff between learnability and memory locality. It solves the linear system equivalent using operations similar to that of the Power method, to that of (PPR-symm), with $\beta=\frac{1-\alpha}{1+\alpha}\in(0,1)$ as the teleportation parameter.
@@
The following Lemma from Andersen et al. (2006) form the basis of the memory locality guarantee of this method. The steps within the while loop are often referred to collectively as the Push operation.
@@ 3 Influencer-targeted sparsification
Figure 2: Is it better to subsample influencers? Correlation of inverse resistive distance (conductance, spectrally optimal) with high degree connecting edge is largely positive, across several graphs.
@@ 3.1 Resistive distance
Figure 3 investigates the performance of edge subsampling as it pertains to node labeling, our desired downstream task. Specifically, it gives the proportion of edges that connect different-labeled nodes over same-labeled nodes. (Smaller is better.) This suggests that influencer-based sparsity not only preserves but can sometimes improve the edge ratio despite a reduction in the number of edges (even moreso than resistive subsampling).
@@ 3.2 Graph sparsification
Algorithm 2 General offline sparsifier 1:Graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},\mathbf{A})$, 2:edge probabilities $p_{u,v}$ for $(u,v)\in{\bm{E}}$ 3:Initialize $\tilde{\mathcal{E}}=\emptyset$, $\tilde{{\bm{A}}}=\mathbf{0}$, $\tilde{\mathcal{G}}=({\mathcal{V}},\tilde{\mathcal{E}},\tilde{{\bm{A}}})$ 4:for $i=1,...,m$ do 5:     Randomly pick edge $(u,v)$. 6:     With probability $p_{u,v}$, 7:$\tilde{{\bm{A}}}_{u,v}=\tilde{{\bm{A}}}_{v,u}=\frac{1}{p_{u,v}},\;\tilde{\mathcal{E}}=\tilde{\mathcal{E}}\bigcup\{(u,v)\}$ return $\tilde{\mathcal{G}}=({\mathcal{V}},\tilde{\mathcal{E}},\tilde{\mathbf{A}})$ [Uncaptioned image] Figure 3: Edge ratio for political-blog. Here, $\bar{q}=c\times$ median distribution for (inf.), and the resulting sparsity level is used for $p$ in (res.) and (unif.).
@@
Consider ${\bm{L}}={\bm{I}}-{\bm{D}}^{-1/2}{\bm{A}}{\bm{D}}^{-1/2}$, $\tilde{\bm{L}}$ the Laplacian matrices corresponding to a graph and its sparsified version, and $\bar{q}$ the subsampling threshold. Then, for any ${\bm{x}}\in\mathbb{R}^{n}$,
$\mathbf{Pr}(|{\bm{x}}^{T}\tilde{\bm{L}}_{I}{\bm{x}}-{\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}|\geq\epsilon)\leq 2\min\{e^{-\frac{\epsilon^{2}}{8\|{\bm{x}}\|_{\infty}^{2}\|{\bm{x}}\|_{2}^{2}}},e^{-\frac{\epsilon^{2}}{8\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|}}\}.$
For a normalized $\|{\bm{x}}\|_{2}=O(1)$, useful values of $\epsilon\in(0,2)$. The bound’s usefulness also depends on the choice of ${\bm{x}}$, and the ratio $\frac{\|x\|_{2}}{\|x\|_{\infty}}$. For a Gaussian random vector in $\mathbb{R}^{n}$, this ratio is about $O(\sqrt{n/\log(n)})$. Plugging in this trend, the right-hand side reduces to $O(\min\{e^{-\frac{\epsilon^{2}n}{8\log(n)}},e^{-\frac{\epsilon^{2}n^{2}}{8\log(n)^{2}|{\mathcal{S}}_{I}|}}\}).$ Note that both terms are nontrivial for very large $n$, with stronger rates if $|{\mathcal{S}}_{I}|<O(n/\log(n))$.
Corollary 3.1.1.
For $n$ large enough,
$\mathbf{Pr}({\bm{L}}_{I}-\epsilon{\bm{I}}\preceq\tilde{\bm{L}}_{I}\preceq{\bm{L}}_{I}+\epsilon{\bm{I}})\geq 1-12\exp(-\frac{\epsilon^{2}n^{3}}{32\log(n)})-4\exp\left(-c^{\prime}n\epsilon/2\right)$
@@ 5.2 Unsupervised clustering
$\mathbf{cluster}(u)=\textsc{PickOne}(\underset{j\in{\mathcal{C}}}{\mathrm{arg\,max}}\{{\bm{Z}}_{t,j}\})\vspace{-2ex}$
@@ 6 Numerical experiments
Figure 6 gives the performance of offline and online (with and without dual correction) methods for online node labeling. Note that for small graphs, the baseline (WMA) is very strong; however, for larger graphs, the tradeoff for reduced misclassification rate becomes more apparent. It is also clear that the Relaxation method, although providing strong rates, is not as strong in practice as the Regularization method. Nonetheless, both methods are closely related, and their respective advantages and disadvantages stem from subtle adjustments in procedures and hyperparameters.
@@ 7 Discussion
The impact of downweighting influencers.
Figures 3 and 10 (appendix) show how different offline subsampling methods affect the edge ratio, which correlates graph clustering features with true labels. Note that influencers and inverse resistive distances are correlated, but not identical; for example, edges between nodes in a fully connected subgraph will have low resistive distance, but are not necessarily influencers. In the political-blog dataset (Fig 3), it seems beneficial to subsample edges from extreme influencers who may ignore political affiliations. However, because low resistive distances might also indicate cliques tied to political parties, indiscriminately removing these can be harmful. These dynamics differ across datasets and applications.
@@ B.1 General facts about subgaussianity
${\mathbf{subgauss}}(X)=S(p):=\begin{cases}1/4&1/4<S_{1}(p),0\leq p\leq 1\\ S_{1}(p)&1/4\geq S_{1}(p),0\leq p\leq 1\\ 1&p>1\\ \end{cases}$
@@ B.2 Offline sparsification
This extra result derives the subgaussian bound for this particular subsampling problem. However, it is not essential for proving our main result for offline sparsification.
@@
${\mathbf{subgauss}}({\bm{x}}^{T}\tilde{{\bm{L}}}_{I}{\bm{x}})=S(\frac{\bar{q}}{d_{\max}})\frac{d_{\max}^{2}}{4\bar{q}^{2}}\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|,\qquad\mathbf{var}({\bm{x}}^{T}\tilde{{\bm{L}}}_{I}{\bm{x}})=\frac{\bar{q}}{\max\{d_{i},d_{j}\}}\cdot(1-\frac{\bar{q}}{\max\{d_{i},d_{j}\}})\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|$
@@
Define $\tilde{{\bm{L}}}_{I}=\tilde{{\bm{L}}}-{\bm{L}}_{C}$, $\bm{L}_{I}-\bm{L}-\bm{L}_{C}$.
@@
and 1 for $(i,j)\not\in{\mathcal{S}}_{I}$. Then $\mathbb{E}[\tilde{{\bm{L}}}]={\bm{L}}$ and ${\bm{x}}^{T}\tilde{{\bm{L}}}{\bm{x}}$ is subgaussian with parameter $\sigma^{2}$ if and only if ${\bm{x}}^{T}\tilde{{\bm{L}}_{I}}{\bm{x}}$ is subgaussian with parameter $\sigma^{2}$. Since $c_{i,j}$ is a scaled Bernoulli random variable which is subgaussian with parameter $\sigma_{i,j}^{2}=\frac{S(p_{i,j})}{p_{i,j}^{2}}=S(\frac{\bar{q}}{\max\{d_{i},d_{j}\}})\frac{\max\{d_{i},d_{j}\}^{2}}{\bar{q}^{2}}$ and
@@
$\sum_{(i,j)\in{\mathcal{S}}_{I}}({\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j})^{2}c_{i,j}^{2}=\sum_{(i,j)\in{\mathcal{S}}_{I}}({\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j})^{2}S(\frac{\bar{q}}{\max\{d_{i},d_{j}\}})\frac{\max\{d_{i},d_{i}\}^{2}}{\bar{q}^{2}}\\ \leq S(\frac{\bar{q}}{\max\{d_{i},d_{j}\}})\frac{\max\{d_{i},d_{j}\}^{2}}{\bar{q}^{2}}\sum_{(i,j)\in{\mathcal{S}}_{I}}({\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j})^{2}.$
Similarly, taking the random variable $Z={\bm{x}}^{T}\bm{L}_{I}{\bm{x}}$, then since $-1\leq\bm{L}_{i,j}\leq 1$, then the following random variable is bounded
$\mathbb{E}[Z]={\bm{x}}^{T}\bm{L}_{I}{\bm{x}},\qquad 0\leq{\bm{x}}^{T}\bm{L}_{I}{\bm{x}}\leq\|{\bm{x}}\|_{\infty}^{2}|{\mathcal{S}}_{I}|.$
and
$\mathbf{var}({\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}})=\sum_{(i,j)\in{\mathcal{S}}_{I}}({\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j})^{2}\mathbf{var}(c_{i,j})\\ =\sum_{(i,j)\in{\mathcal{S}}_{I}}({\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j})^{2}\frac{\bar{q}}{\max\{d_{i},d_{j}\}}\cdot(1-\frac{\bar{q}}{\max\{d_{i},d_{j}\}})\leq\frac{1}{4}\sum_{(i,j)\in{\mathcal{S}}_{I}}({\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j})^{2}$
@@
${\mathbf{subgauss}}({\bm{x}}^{T}\tilde{{\bm{L}}}_{I}{\bm{x}})=S(\frac{\bar{q}}{d_{\max}})\frac{d_{\max}^{2}}{\bar{q}^{2}}\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|,\qquad\mathbf{var}({\bm{x}}^{T}\tilde{{\bm{L}}}_{I}{\bm{x}})=\frac{\|{\bm{x}}\|_{\infty}^{4}}{4}|{\mathcal{S}}_{I}|$
Theorem B.5 (Offline sparsification).
The following holds pointwise over all ${\bm{x}}$
$\mathbf{Pr}(|{\bm{x}}^{T}\tilde{\bm{L}}_{I}{\bm{x}}-{\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}|\geq\epsilon)\leq 2\min\{e^{-\frac{\epsilon^{2}}{8\|{\bm{x}}\|_{\infty}^{2}\|{\bm{x}}\|_{2}^{2}}},e^{-\frac{\epsilon^{2}}{8\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|}}\}.$
Proof.
The first two results are direct applications of Bernstein’s inequality. The last result is proven below.
Suppose $Z\sim\mathbf{Bern}(0,R)$ with mean $\mu$. Then for all $0\leq\theta\leq 1$,
[…]
Then, since $|{\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j}|\leq\|{\bm{x}}\|_{\infty}^{2}$,
$\mathbb{E}[e^{\lambda{\bm{x}}^{T}\tilde{\bm{L}}_{I}{\bm{x}}}]=\mathbb{E}[e^{\lambda\sum_{(i,j)\in{\mathcal{S}}_{I}}{\bm{x}}_{i}{\bm{x}}_{j}(\tilde{\bm{L}}_{I})_{i,j}}]\\ \leq e^{\lambda\sum_{(i,j)\in{\mathcal{S}}_{I}}{\bm{x}}_{i}{\bm{x}}_{j}{\bm{L}}_{i,j}(1+\lambda\|{\bm{x}}\|_{\infty}^{2})}=e^{\lambda{\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}(1+\lambda\|{\bm{x}}\|_{\infty}^{2})},\qquad\lambda\|{\bm{x}}\|_{\infty}^{2}\leq 1.$
By Chernoff’s inequality,
$\mathbf{Pr}({\bm{x}}^{T}\tilde{\bm{L}}_{I}{\bm{x}}\geq(1+\epsilon){\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}})$$\leq$ $\frac{\mathbb{E}[e^{\lambda{\bm{x}}^{T}\tilde{\bm{L}}_{I}{\bm{x}}}]}{e^{\lambda(1+\epsilon){\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}}}\quad\forall\lambda>0$
$\leq$ $e^{\lambda{\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}(\lambda\|{\bm{x}}\|_{\infty}^{2}-\epsilon)}\quad\forall 0<\lambda\leq\frac{1}{\|{\bm{x}}\|_{\infty}^{2}}$
$=$ $e^{-\frac{{\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}}{4\|{\bm{x}}\|_{\infty}^{2}}\epsilon^{2}},\quad\lambda=\frac{\epsilon}{2\|{\bm{x}}\|_{\infty}^{2}}$
Extending to a two-sided bound and using $\tilde{\epsilon}=\epsilon{\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}$,
$\mathbf{Pr}(|{\bm{x}}^{T}\tilde{\bm{L}}_{I}{\bm{x}}-{\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}|\geq\tilde{\epsilon})\leq 2e^{-\frac{\tilde{\epsilon}^{2}}{4\|{\bm{x}}\|_{\infty}^{2}{\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}}}\leq 2\min\{e^{-\frac{\epsilon^{2}}{8\|{\bm{x}}\|_{\infty}^{2}\|{\bm{x}}\|_{2}^{2}}},e^{-\frac{\epsilon^{2}}{8\|{\bm{x}}\|_{\infty}^{4}|{\mathcal{S}}_{I}|}}\}.$
since in the last inequality, ${\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}\leq\|{\bm{x}}\|_{2}^{2}\|{\bm{L}}_{I}\|_{2}\leq\|{\bm{x}}\|_{2}^{2}$ or also ${\bm{x}}^{T}{\bm{L}}_{I}{\bm{x}}\leq\|{\bm{x}}\|_{\infty}^{2}|{\mathcal{S}}_{I}|$.
Lemma B.6.
For ${\bm{x}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{n})$, for $n>3$,
$\mathbf{Pr}(\frac{\|{\bm{x}}\|_{\infty}}{\|{\bm{x}}\|_{2}}\geq\sqrt{\frac{\log(n)}{n}})\leq(n+1)\exp(-\frac{n\log(n)}{2})$
Proof.
From Vershynin (2018), for a Gaussian random variable $x$,
$\mathbf{Pr}(x\geq\epsilon)\leq\frac{1}{\sqrt{2\pi}}\exp(-\frac{\epsilon^{2}}{2}),\qquad\epsilon\geq 1.$
Applying union bound,
$\mathbf{Pr}\left(\|{\bm{x}}\|_{\infty}\geq t_{1}\right)\leq\frac{n}{\sqrt{2\pi}}\exp(-\frac{t_{1}^{2}}{2}),\qquad t_{1}\geq 1.$
Also from Vershynin (2018),
$\mathbf{Pr}(|\|x\|_{2}^{2}-n|\geq\epsilon)\leq 2\exp(-\frac{1}{2}\min(\frac{\epsilon^{2}}{n},\frac{\epsilon}{2}))$
so picking $t_{2}^{2}=n+\epsilon$,
$\Pr\left(\|{\bm{x}}\|_{2}^{2}\leq t_{2}^{2}\right)\leq 2\exp(-\frac{(n-t_{2}^{2})^{2}}{2n})$
Hence,
$\mathbf{Pr}(\frac{\|{\bm{x}}\|_{\infty}}{\|{\bm{x}}\|_{2}}\leq\frac{t_{1}}{t_{2}})\geq\mathbf{Pr}(\|{\bm{x}}\|_{\infty}\leq t_{1}\text{ and }\|{\bm{x}}\|_{2}\geq t_{2})\\ =1-\mathbf{Pr}(\|{\bm{x}}\|_{\infty}\geq t_{1}\text{ or }\|{\bm{x}}\|_{2}\leq t_{2})\geq 1-\frac{n}{\sqrt{2\pi}}\exp(-\frac{t_{1}^{2}}{2})-2\exp(-\frac{(n-t^{2}_{2})^{2}}{2n})$
so,
$\mathbf{Pr}(\frac{\|{\bm{x}}\|_{\infty}}{\|{\bm{x}}\|_{2}}\geq\frac{t_{1}}{t_{2}})\leq\frac{n}{\sqrt{2\pi}}\exp(-\frac{t_{1}^{2}}{2})+2\exp(-\frac{(n-t^{2}_{2})^{2}}{2n}).$
Taking $t_{2}=n$ and $t_{1}^{2}=n\log(n)$ yields
$\mathbf{Pr}(\frac{\|{\bm{x}}\|_{\infty}}{\|{\bm{x}}\|_{2}}\geq\sqrt{\frac{\log(n)}{n}})\leq\frac{n}{\sqrt{2\pi}}\exp(-\frac{n\log(n)}{2})+2\exp(-\frac{(n-n^{2})^{2}}{2n})\leq(n+1)\exp(-\frac{n\log(n)}{2})$
for $n>3$. ∎
Lemma B.7 ( Rudelson & Vershynin (2013) Thm. 2.1).
Let ${\bm{x}}_{1},\dots,{\bm{x}}_{k}\sim\mathcal{N}(0,{\bm{I}}_{n})$ be i.i.d. Gaussian vectors and ${\bm{A}}\in\mathbb{R}^{n\times n}$. Then
$\mathbf{Pr}(|\|{\bm{A}}{\bm{x}}\|_{2}-\|{\bm{A}}\|_{F}|>\epsilon)\leq 2\exp(-\frac{c\epsilon^{2}}{s^{4}\|{\bm{A}}\|^{2}_{2}})$
where $c$ is a constant that does not depend on ${\bm{A}}$ or $n$, and $s$ is the subgaussian constant of a Gaussian random variable.
This is a consequence of the Hanson-Wright inequality.